Find area of curve y^2=2x and line y=4x
Answer by Alan3354(69443)

Find area of curve y^2=2x and line y=4x
--------------
Find area of curve x = y^2/2 and line x=y/4
----
The 2 points of intersection are (0,0) and (1/8,1/2)
Integrate wrt to y
x = y^2/2
INT = y^3/6
---
x=y/4
INT = y^2/8
------------
(Ignore the constants of integration)
Area = y^3/6 - y^2/8 from y = 0 to y = 1/2
Area = (1/2)^3/6 - (1/2)^2/8
= 1/48 - 1/32
= 1/96 sq units